OpenStudy (anonymous):

What is the solubility of AgBr in g/L in 0.0010 M AgNO3 solution? The Ksp for AgBr is 7.7 x 10−13. A. 1.5 x 10−7 g/L B. 7.7 x 10−11 g/L C. 3.6 x 10−7 g/L D. 2.2 x 10−5 g/L E. 4.2 x 10−10 g/L

OpenStudy (jfraser):

KSP lets us find the maximum solubility of a solute in terms of Mol/L, so we can convert from mol/L into g/L by using the molar mass. Use the KSP equation\[K_{SP} = [Ag{^+}][Br{^-}] = 7.7*10^{-13}\] to solve for the concentration of Ag+ ions in solution. Since the ratio of ions in AgBr is 1:1, whatever the concentration of silver ions, the bromide will be the same.\[7.7*10^{-13} = [x][x] = x^2\] That concentration is in mol/L, so use the molar mass of AgBr to turn those moles of AgBr into grams of AgBr.